# Estimating the Reynolds number dependence of the chaotic attractor dimension in two-dimensional Kolmogorov flow

### v1

@@ I Introduction
Nonlinear partial differential equations (PDEs), such as the Navier–Stokes equations that govern turbulent fluid motion, are fundamental in modeling complex dynamics and often exhibit spatiotemporal chaos in parameter regimes of practical interest. Although the state spaces of PDEs are formally infinite-dimensional, it is argued that solutions of the Navier-Stokes equations collapse onto a finite-dimensional inertial manifold, denoted here by $\mathcal{M}$, due to the dissipative action of viscosity [1, 2, 3, 4]. Although the existence of inertial manifolds for the Navier-Stokes equations has not been proven [5], approximate inertial manifolds, which attract all orbits within a thin neighborhood of the global attractor, have been constructed in two spatial dimensions [6]. The inertial manifold encapsulates the essential features of the system, with an intrinsic dimension $d_{\mathcal{M}}$, typically much smaller than the dimension $d_{\omega}$ of the state space of the discretized PDE.
[…]
@@
Our focus in this work is on generating low-dimensional embeddings for the chaotic attractor $\mathcal{A}\subset\mathcal{A}_{G}$ in two-dimensional Kolmogorov flow, which describes the long-time evolution of turbulence and which presumably requires far fewer degrees of freedom than $\mathcal{A}_{G}$. To estimate the scaling $d_{\mathcal{A}}:=\dim(\mathcal{A})$ with Reynolds number, $Re$, we construct highly-accurate autoencoder neural networks in an effort to find robust ‘latent’ coordinate systems. Our objectives are twofold: (1) to estimate the $Re$-dependence of $d_{\mathcal{A}}$ and (2) to use the latent coordinate system to map typical dynamical pathways through different vortical ‘events’ and to understand the coverage (or lack of coverage) of these dynamics by large libraries of recently computed unstable periodic orbits [13, 14].
@@
Numerical computations for the chaotic attractor dimension (by computing Lyapunov exponents and invoking the Kaplan-Yorke conjecture [15]) give some indications as to a scaling with $Re$. The Lyapunov dimension of the attractor of Kolmogorov flow was reported to depend strongly on the wavenumber $n$ chosen for the monochromatic forcing profile, and only weakly on $Re\propto G^{1/2}$ [16]. A follow-on study of Kolmogorov flow found that the Lyapunov dimension is bounded above by the number of degrees of freedom contained in scales larger than the forcing scale, $N_{LS}\propto n^{2}$ [17], though this study considered relatively weak turbulence with $16.4\leq Re\leq 46.2$ in our units. A scaling of the Lyapunov dimension $\sim Re^{0.78}(n/k_{0})^{2}$ was also computed numerically for forced two-dimensional turbulence with large scale damping, where $k_{0}=2\pi/L$ [18], which also has a linear dependence on $N_{LS}$.
@@
Many canonical fluid flows exhibit continuous and/or discrete symmetries in the governing equations, which must be accounted for in the training of a nonlinear neural network. The simplest way to encourage the model to learn the symmetries of the equations is via data augmentation, where each symmetry is randomly applied to the training data after each epoch [13, 30] to ensure a robust sampling of the symmetry groups. However, this approach does not guarantee equivariance of the model with respect to the symmetries. In this work, to explicitly reduce a continuous or discrete symmetry, each snapshot is mapped a symmetry-reduced subspace of the full state space. This pre-processing of the data restricts the model to learning a representation within this subspace [31, 28]. The first Fourier mode method of slices can be used to explicitly reduce continuous symmetries of dynamical systems [32]. For discrete symmetries, we adapt a reduction technique that defines an indicator function based on Fourier modes to separate the state space of the governing equation into symmetric charts, and maps all solutions into one of these charts, defining a ‘fundamental chart’ [this was applied to the KS system in 31]. This method of ‘symmetry charting’ has been applied to Kolmogorov flow with forcing wavenumber $n=2$ [28], where an indicator function was defined on specific Fourier modes to define eight symmetry charts representing the eight-fold discrete symmetries. Other work has explicitly accommodated various discrete symmetries via the use of many identical (‘Siamese’) networks to identify a fundamental domain [33].
@@ II.1 Numerical simulations
[…]
@@ II.2.1 Resampling
[…]
@@ II.2.2 Symmetry reduction
As described in §II.1, Kolmogorov flow is equivariant under continuous shifts in the streamwise direction, under discrete shift-reflects by a half-wavelength in $y$, and under rotations by $\pi$. We account for all symmetries explicitly in a pre-processing step before training the neural networks The continuous translational symmetry is reduced by the first Fourier mode method of slicing [32]. In this method a phase $\phi$ is computed such that the $(k_{x},k_{y})=(1,0)$ Fourier mode of the translated state $\mathscr{T}_{-\phi}\omega$ is purely real.
@@
We consider the $(0,1)$ mode to diagnose the shift-reflect symmetry, splitting the complex plane spanned by the real $\hat{\omega}_{R}(0,1)$ and imaginary $\hat{\omega}_{I}(0,1)$ components of this mode into octants, representing the 8 shift-reflected copies generated by (6). We consider the $(0,4)$ mode for the rotational symmetry, splitting this complex plane into upper and lower halves, i.e. defined by the sign of $\hat{\omega}_{I}(0,4)$. Together, this decomposition results in 16 indicators as indicated in figure 2, representing the 16 symmetric copies generated by both (6) and (7). To reduce the discrete symmetries of a state, we select the ‘fundamental’ domain as $\mathscr{I}=1$. The required combination of symmetry operations to map a state from any $\mathscr{I}$ to $\mathscr{I}=1$ is tabulated in table 1.
Generalization of this approach to any number of forcing waves $n$ is relatively straightforward – it requires the partitioning of the complex plane spanned by $\hat{\omega}_{R}(0,1)$ and $\hat{\omega}_{I}(0,1)$ into $n$ equally-sized sectors, along with consideration of the sign of $\hat{\omega}_{I}(0,n)$ to diagnose rotation.
@@ 16 $\mathscr{R}$
Table 1: Symmetry operations to map all 16 symmetry charts of $n=4$ Kolmogorov flow to the fundamental chart $\mathcal{I}=1$
@@ III Network architecture
We construct low-rank representations of the turbulent attractor in Kolmogorov flow by combining the DenseNet convolutional autoencoder architecture from a previous study [13] (hereafter PHBK) with the implicit rank minimization and weight decay introduced by the IRMAE-WD architecture [26, 28]. Each deep convolutional autoencoder $\mathscr{A}$ seeks to reconstruct its input such that
@@
where the encoder $\mathscr{E}:\mathbb{R}^{N_{x}\times N_{y}}\to\mathbb{R}^{d_{z}}$ maps the input vorticity snapshot to a low-dimensional representation, $\mathscr{W}:\mathbb{R}^{d_{z}}\to\mathbb{R}^{d_{z}}$ represents a series of $n$ fully-connected, equally-sized linear layers ( pure matrix multiplication) within the embedding space such that $\mathscr{W}=\mathbf{W}_{1}\mathbf{W}_{2}\cdots\mathbf{W}_{n}$ and $\mathbf{W}_{i}\in\mathbb{R}^{d_{z}\times d_{z}}$, and the decoder $\mathscr{D}:\mathbb{R}^{d_{z}}\to\mathbb{R}^{N_{x}\times N_{y}}$ maps the embedding back to a vorticity snapshot.
The structures of the encoder $\mathscr{E}$ and decoder $\mathscr{D}$ considered here are adapted from PHBK, with so-called ‘dense blocks’ [42, 43] instead of single convolutional layers. The dense blocks allow for much richer feature maps at each scale throughout the autoencoder. Full details on the architecture used here are given in appendix A. The key addition to the architecture is the insertion of fully-connected linear layers into the bottleneck of the autoencoder. The implicit rank minimization of these linear layers drives the output of $\mathscr{W}(\mathscr{E}(\omega))$ to be of minimal rank [27]. The densely connected linear layers to the autoencoder bottleneck do not increase the expressivity of the network, as these layers can be expressed as a multiplication by a square matrix. However, the number of parameters increases to $\sim 6.21\times 10^{6}$ trainable parameters due to the insertion of these linear layers, nearly triple the $\sim 2.15\times 10^{6}$ trainable parameters in the comparable models considered in PHBK. The ‘IRMAE’ approach – while fully deterministic – has been shown to result in latent (or embedding) spaces with many favorable properties reminiscent of more complex approaches (e.g. variational autoencoders), including ability to smoothly interpolate between embeddings and to generate new samples [27].
@@
using an AdamW optimizer [44] with an initial learning rate of $10^{-4}$ and the default weight decay coefficient of 0.004. This optimizer decouples the weight decay from the adaptive gradient updates, as including an explicit $L_{2}$ regularization term on the trainable parameters in (9) is not effective with Adam-style optimizers [45], nor identical to weight decay [44]. The batch size is set to 8 for $Re\leq 40$, 12 for $Re=60$ and 16 for $Re\geq 80$. Setting relatively small batch size increases the total training time, but also increases the total number of gradient-based optimization steps which we found to to be effective in driving down the rank of the learned embedding. The networks are trained with the AdamW optimizer for at least 1000 epochs for each $Re$. Inspired by previous studies implementing a two-optimizer training process [46], the best model to that point is then further tuned using the AdaGrad optimizer [47] for a further 1000 epochs.
@@ IV.1 Reconstruction error
Figure 3: The average reconstruction error $\varepsilon$ of the model against $Re$. The error bars at each $Re$ are defined by one standard deviation of the reconstruction error.
@@
The distribution of the error and correlation with dissipation rate is examined for a sample of the $Re$values considered in figure 5. For the $Re=40$ results, we also compare the performance of identical architectures trained on dissipation-resampled (color) or not resampled (gray) data at $Re=40$. For a fair comparison, a similar number of training snapshots were used in both cases. The resampling strategy considerably improves the error of the high dissipation events, and surprisingly also slightly improves the error of the low dissipation events. The reconstruction error is still slightly correlated with dissipation rate, especially at higher $Re$.
@@
We report the test-averaged energy spectrum $R(k)$ of the residual fields $r_{j}=\mathcal{A}(\omega_{j})-\omega_{j}$ generated by the model at $Re=40$ and $Re=100$ in figure 6. We have replaced the traditional single output layer in our architecture by a sequence of linear convolutional layers of decreasing kernel sizes from $(16,16)$ to $(6,6)$ (full details in appendix A). This ‘output stream’ of layers was motivated by the empirical observation that the reconstructed energy spectrum with a single output convolutional layer plateaued incorrectly at high wavenumbers. This plateau is evident in figure 6, where the red residual spectrum is generated by an identical fully-trained architecture with a single linear output layer of filter size $(8,8)$ instead of the output stream. Empirically, the reduction in filter size through the stream tends to sequentially refine at finer spatial scales. Comparing the residual spectra of the output with and without the output stream at $Re=40$, we see a notable improvement at the largest spatial scale, and a slight improvement at the very small spatial scales. We found that using any smaller filter sizes resulted in an over-smoothing of the reconstructed vorticity fields.
Figure 6: The residual spectrum $R(k)$ generated by the model shown in blue at (left) $Re=40$ and (right) $Re=100$. The red line indicates the residual spectrum at $Re=40$ of a model trained with a single linear convolutional output layer of filter size $(8,8)$.

### v2

# Characterizing the Reynolds number dependence of the chaotic attractor in two-dimensional turbulence with dimension-minimizing autoencoders

@@ I Introduction
Nonlinear partial differential equations (PDEs), such as the Navier–Stokes equations that govern turbulent fluid motion, are fundamental in modeling complex dynamics and often exhibit spatiotemporal chaos in parameter regimes of practical interest. Although the state spaces of PDEs are formally infinite-dimensional, it is argued that solutions of the Navier-Stokes equations collapse onto a finite-dimensional inertial manifold, denoted here by $\mathcal{M}$, due to the dissipative action of viscosity [1, 2, 3, 4]. The inertial manifold is a smooth, finite-dimensional manifold, within which the global attractor (which may be a complicated fractal set) is embedded [3, 4]. The global attractor $\mathcal{A}_{G}$ is defined as the set of points in phase space that can be reached in arbitrarily long time from all initial conditions [5]. Two familiar properties of $\mathcal{A}_{G}$ are (i) it attracts the time forward map of all initial conditions $\bm{u}_{0}$, such that $\text{dist}(\bm{f}^{t}(\bm{u}_{0}),\mathcal{A}_{G})\to 0$ as $t\to\infty$, and (ii) it is an invariant set. Although the existence of inertial manifolds for the Navier-Stokes equations has not been proven [6], approximate inertial manifolds, which attract all orbits within a thin neighborhood of the global attractor, have been constructed in two spatial dimensions [7]. The inertial manifold encapsulates the essential features of the system, with an intrinsic dimension $d_{\mathcal{M}}$, typically much smaller than the dimension $d_{\omega}$ of the state space of the discretized PDE.
While the existence of an inertial manifold is not proven, the global attractor $\mathcal{A}_{G}\subset\mathcal{M}$ exists for the forced 2D Navier-Stokes equations [2], and its dimension has been bounded from above [8, 9] and below [10] (the latter in monochromatic Kolmogorov flow). The bounds are asymptotically tight (up to a logarithmic factor): The upper bound scales with a non-dimensional Grashof number $G:=(\int_{A}|\mathbf{f}|^{2}dA)^{1/2}\,L^{2}/\nu^{2}$ as $G^{2/3}(1+\log G)^{1/3}$ [8, 9], which depends on the forcing profile $\mathbf{f}$, box length $L$ and viscosity $\nu$, while the lower bound scales like $G^{2/3}$ [10] (this computation is based on the dimension of the unstable manifold of the laminar fixed point). Alternatively, a heuristic argument for the dimension of the inertial manifold has been made by determining the total number of excited modes in the inertial range of the (direct) turbulent cascade. This estimate scales like $N\sim(L/l_{d})^{2}\propto G^{2/3}$ in two-dimensions, where $L$ is the largest length scale in the domain and $l_{d}$ is the dissipative (Kraichnan) length scale [11]. This heuristic estimate was extended to $G^{2/3}(\log G)^{1/3}$ by incorporating a logarithmic correction to the energy spectrum to ensure a constant enstrophy flux in the direct cascade [12], and was shown to be consistent with the rigorous upper bound [13].
@@
Our focus in this work is on generating low-dimensional embeddings for the chaotic attractor $\mathcal{A}\subset\mathcal{A}_{G}$ in two-dimensional Kolmogorov flow. $\mathcal{A}$ is defined as an invariant set which attracts an open set of initial conditions and is minimal – no proper subset of $\mathcal{A}$ exists which satisfies the first two conditions. The chaotic attractor presumably requires far fewer degrees of freedom than $\mathcal{A}_{G}$, which must contain all exact coherent structures and their unstable manifolds. To estimate the scaling $d_{\mathcal{A}}:=\dim(\mathcal{A})$ with Reynolds number, $Re$, we construct highly-accurate autoencoder neural networks in an effort to find robust ‘latent’ coordinate systems. Our objectives are twofold: (1) to estimate the $Re$-dependence of $d_{\mathcal{A}}$ and (2) to use the latent coordinate system to map typical dynamical pathways through different vortical ‘events’ and to understand the coverage (or lack of coverage) of these dynamics by large libraries of recently computed unstable periodic orbits [14, 15].
@@
Numerical computations for the chaotic attractor dimension (by computing Lyapunov exponents and invoking the Kaplan-Yorke conjecture [16]) give some indications as to a scaling with $Re$. The Lyapunov dimension of the attractor of Kolmogorov flow was reported to depend strongly on the wavenumber $n$ chosen for the monochromatic forcing profile, and only weakly on $Re\propto G^{1/2}$ [17]. A follow-on study of Kolmogorov flow found that the Lyapunov dimension is bounded above by the number of degrees of freedom contained in scales larger than the forcing scale, $N_{LS}\propto n^{2}$ [18], though this study considered relatively weak turbulence with $16.4\leq Re\leq 46.2$ in our units. A scaling of the Lyapunov dimension $\sim Re^{0.78}(n/k_{0})^{2}$ was also computed numerically for forced two-dimensional turbulence with large scale damping, where $k_{0}=2\pi/L$ [19], also demonstrating a linear dependence on $N_{LS}$.
@@
Many canonical fluid flows exhibit continuous and/or discrete symmetries in the governing equations, which must be accounted for in the training of a nonlinear neural network. The simplest way to encourage the model to learn the symmetries of the equations is via data augmentation, where each symmetry is randomly applied to the training data after each epoch [14, 31] to ensure a robust sampling of the symmetry groups. However, this approach does not guarantee equivariance of the model with respect to the symmetries. In this work, to explicitly reduce a continuous or discrete symmetry, each snapshot is mapped a symmetry-reduced subspace of the full state space. This pre-processing of the data restricts the model to learning a representation within this subspace [32, 29]. The first Fourier mode method of slices can be used to explicitly reduce continuous symmetries of dynamical systems [33]. For discrete symmetries, we adapt a reduction technique that defines an indicator function based on Fourier modes to separate the state space of the governing equation into symmetric charts, and maps all solutions into one of these charts, defining a ‘fundamental chart’ [this was applied to the KS system in 32]. This method of ‘symmetry charting’ has been applied to Kolmogorov flow with forcing wavenumber $n=2$ [29], where an indicator function was defined on specific Fourier modes to define eight symmetry charts representing the eight-fold discrete symmetries. Other work has explicitly accommodated various discrete symmetries via the use of many identical (‘Siamese’) networks to identify a fundamental domain [34].
@@ II.1 Numerical simulations
We consider this flow at a number of $Re\in\{$25, 30, 35, 40, 60, 80, 100, 200, 400}\}. Kolmogorov flow approaches an asymptotic regime [38, 39] with $D\sim Re^{-1/2}$ beyond $Re\approx 50$. A variety of simpler states have been reported at lower $Re$ [38]. The majority of $Re$ values considered are sampled from the asymptotic chaotic regime. We will devote particular attention to the well-studied values $Re=40$ and $Re=100$.
@@ II.2.1 Resampling
Training models on data sampled from the turbulent attractor, as described above, lead to low reconstruction errors for low-dissipation-rate snapshots, but high errors for high-dissipation-rate snapshots as observed in previous studies [35, 14]. This is because (1) low-dissipation snapshots are less spatially complex and (2) high-dissipation events are rare events in Kolmogorov flow [38, 42] – the dataset is dominated by low-dissipation events (see figure 1). One approach that has previously been used to alleviate this effect is the introduction of a loss term involving the square of the vorticity field [14] to add additional penalization to higher dissipation events. Here, we instead introduce a resampling strategy to alter the distribution of training data over dissipation and artificially increase the frequency of high dissipation events. To do this, the distribution of dissipation rate is first partitioned into ‘low’ and ‘high’ dissipation events, defined by a threshold value $D_{c}$. For example, at $Re=100$, this value is set to $D_{c}=0.045D_{lam}$ (the laminar dissipation rate is $D_{lam}=Re/2n^{2}$). All data above the threshold value is sampled from the original training dataset. The dataset is then binned according to dissipation rate, with bin widths of $\Delta D=0.001D_{lam}$ for $Re<200$, $\Delta D=0.0005D_{lam}$ for $Re=200$ and $\Delta D=0.0001D_{lam}$ for $Re=400$. To smoothly reduce the number of low dissipation events in the dataset, a maximum of $n_{l}$ snapshots are then randomly sampled from each bin below the threshold $D_{c}$, where $n_{l}$ is set by the number of samples in the first ‘high’ dissipation bin $D_{c},D_{c}+\Delta D]$. The modified distribution created by this resampling strategy is shown in figure 1 for $Re=100$.
The parameter $D_{c}$ must balance two key considerations: that (1) the network is exposed to a relatively high number of high-dissipation snapshots and that (2) a sufficiently large training dataset is retained after resampling from the $4\times 10^{5}$ original snapshots at each $Re$. These are competing conditions as the size of the resampled dataset can be increased by lowering $D_{c}$, but with a reduced weighting given to the high dissipation snapshots. The thresholds, in order of increasing $Re$, are set at $D_{c}/D_{lam}\in\{$0.16, 0.1415, 0.1255, 0.15, 0.125, 0.0725, 0.045, 0.0125, 0.0045}\}. As a result of the resampling, the size of the training dataset is reduced, with a total number of snapshots snapshots $N_{S}\approx(20,40,50,60,75,90,100,125,130)\times 10^{3}$ (ordered in increasing $Re$).
@@ II.2.2 Symmetry reduction
As described in §II.1, Kolmogorov flow is equivariant under continuous shifts in the streamwise direction, under discrete shift-reflects by a half-wavelength in $y$, and under rotations by $\pi$. We account for all symmetries explicitly in a pre-processing step before training the neural networks. The continuous translational symmetry is reduced by the first Fourier mode method of slicing [33]. In this method a phase $\phi$ is computed such that the $(k_{x},k_{y})=(1,0)$ Fourier mode of the translated state $\mathscr{T}_{-\phi}\omega$ is purely real.
@@
We consider the $(0,1)$ mode to diagnose the shift-reflect symmetry, splitting the complex plane spanned by the real $\hat{\omega}_{R}(0,1)$ and imaginary $\hat{\omega}_{I}(0,1)$ components of this mode into octants, representing the 8 shift-reflected copies generated by (6). We consider the $(0,4)$ mode for the rotational symmetry, splitting this complex plane into upper and lower halves, i.e. defined by the sign of $\hat{\omega}_{I}(0,4)$. Together, this decomposition results in 16 indicators as indicated in figure 2, representing the 16 symmetric copies generated by both (6) and (7). To reduce the discrete symmetries of a state, we select the ‘fundamental’ domain as $\mathscr{I}=1$, which is defined by $0<\arg\hat{\omega}(0,1)<\pi/4$ and $\hat{\omega}_{I}(0,4)>0$.
The required combination of symmetry operations to map a state from any $\mathscr{I}$ to $\mathscr{I}=1$ is tabulated in table 1. In practice, the symmetry reduction of a snapshot proceeds as follows: (i) identify the domain $\mathscr{I}$ of the snapshot from the $(0,1)$ and $(0,4)$ modes as in figure 2, (ii) apply the appropriate symmetry operations from table 1 to the snapshot, (iii) reduce the translational symmetry with the first Fourier mode method of slicing. Generalization of this approach to any number of forcing waves $n$ is relatively straightforward – it requires the partitioning of the complex plane spanned by $\hat{\omega}_{R}(0,1)$ and $\hat{\omega}_{I}(0,1)$ into $n$ equally-sized sectors, along with consideration of the sign of $\hat{\omega}_{I}(0,n)$ to diagnose rotation.
@@ 16 $\mathscr{R}$
Table 1: Symmetry operations to map all 16 symmetry charts of $n=4$ Kolmogorov flow to the fundamental chart $\mathscr{I}=1$
@@ III Network architecture
We construct low- dimensional representations of the turbulent attractor in Kolmogorov flow by combining the DenseNet convolutional autoencoder architecture from a previous study [14] (hereafter PHBK) with the implicit rank minimization and weight decay introduced by the IRMAE-WD architecture [27, 29]. Each deep convolutional autoencoder $\mathscr{A}$ seeks to reconstruct its input such that
@@
where the encoder $\mathscr{E}:\mathbb{R}^{N_{x}\times N_{y}}\to\mathbb{R}^{d_{z}}$ maps the input vorticity snapshot to a low-dimensional representation, $\mathscr{W}:\mathbb{R}^{d_{z}}\to\mathbb{R}^{d_{z}}$ represents a series of $n$ fully-connected, equally-sized linear layers (pure matrix multiplication) within the embedding space such that $\mathscr{W}=\mathbf{W}_{1}\mathbf{W}_{2}\cdots\mathbf{W}_{n}$ and $\mathbf{W}_{i}\in\mathbb{R}^{d_{z}\times d_{z}}$, and the decoder $\mathscr{D}:\mathbb{R}^{d_{z}}\to\mathbb{R}^{N_{x}\times N_{y}}$ maps the embedding back to a vorticity snapshot. A high-level schematic of this architecture and the symmetry reduction pre-processing step is presented in figure 3.
Figure 3: Schematic of the data/neural network pipeline used in this study. Symmetries are first reduced for each snapshot $\omega$, before being passed to the IRMAE network. The implicit rank minimizing layers (orange) are linear fully connected layers, while the encoder $\mathcal{E}$ and decoder $\mathcal{D}$ are fully described in Appendix A.
The structures of the encoder $\mathscr{E}$ and decoder $\mathscr{D}$ considered here are adapted from PHBK, with so-called ‘dense blocks’ [43, 44] instead of single convolutional layers. The dense blocks allow for much richer feature maps at each scale throughout the autoencoder. Full details on the architecture used here are given in appendix A and the full source code is available at https://github.com/computational-flow-science/IRMAE. The key addition to the architecture is the insertion of fully-connected linear layers into the bottleneck of the autoencoder. The implicit rank minimization of these linear layers drives the output of $\mathscr{W}(\mathscr{E}(\omega))$ to be of minimal rank [28]. The densely connected linear layers to the autoencoder bottleneck do not increase the expressivity of the network, as these layers can be expressed as a multiplication by a square matrix. However, the number of parameters increases to $\sim 6.21\times 10^{6}$ trainable parameters due to the insertion of these linear layers, nearly triple the $\sim 2.15\times 10^{6}$ trainable parameters in the comparable models considered in PHBK.
Although more complex (generative) approaches have been shown to learn informative latent representations of chaotic dynamics [45, 46], the IRMAE network – while fully deterministic – has been shown to result in latent (or embedding) spaces with similar favorable properties. For example, comparable performance to variational autoencoders has been demonstrated in tasks such as smoothly interpolating in the latent space and generating new samples from random noise [28]. Furthermore, while the latent spaces of variational autoencoders are typically full rank, the IRMAE networks benefit from approximately minimal rank latent spaces, which are key in estimating the dimensionality of the chaotic attractor.
@@
where each $\omega_{j}$ is a snapshot from the training dataset, using an AdamW optimizer [47] with an initial learning rate of $10^{-4}$ and the default weight decay coefficient of 0.004. This optimizer decouples the weight decay from the adaptive gradient updates, as including an explicit $L_{2}$ regularization term on the trainable parameters in (9) is not effective with Adam-style optimizers [48], nor identical to weight decay [47]. The batch size is set to 8 for $Re\leq 40$, 12 for $Re=60$ and 16 for $Re\geq 80$. Setting relatively small batch size increases the total training time, but also increases the total number of gradient-based optimization steps which we found to be effective in driving down the rank of the learned embedding. The networks are trained with the AdamW optimizer for at least 1000 epochs for each $Re$. Inspired by previous studies implementing a two-optimizer training process [49], the best model to that point is then further tuned using the AdaGrad optimizer [50] for a further 1000 epochs. A brief hyperparameter analysis over the initial learning rate and batch size is presented in appendix C. The model size and associated training time (2 weeks of GPU time per model reported here) prohibited us from performing an extensive hyperparameter sweep over all $Re$, and better performance may well be possible with alternative settings.
@@ IV.1 Reconstruction error
Figure 4: The average reconstruction error $\varepsilon$ of the model against ReRe. A bottleneck dimension of $d_{z}=1024$ is chosen for each $Re$. The error bars at each $Re$ are defined by one standard deviation of the reconstruction error.
@@
The distribution of the error and correlation with dissipation rate is examined for a sample of the $Re$-values considered in figure 6. For the $Re=40$ results, we also compare the performance of identical architectures trained on dissipation-resampled (color) or not resampled (gray) data at $Re=40$. For a fair comparison, a similar number of training snapshots were used in both cases. The resampling strategy considerably improves the error of the high dissipation events, and surprisingly also slightly improves the error of the low dissipation events. The reconstruction error is still slightly correlated with dissipation rate, especially at higher $Re$.
@@
We report the test-averaged enstrophy spectra $\Omega(k)$ generated by the model at $Re=40,100,200$ and $400$ in figure 7. We have replaced the traditional single output layer in our architecture by a sequence of linear convolutional layers of decreasing kernel sizes from $(16,16)$ to $(6,6)$ (full details in appendix A). This ‘output stream’ of layers was motivated by the empirical observation that the reconstructed enstrophy spectrum with a single output convolutional layer plateaued incorrectly at high wavenumbers. This plateau is evident in the top left panel of figure 7, where the red residual spectrum is generated by an identical fully-trained architecture with a single linear output layer of filter size $(8,8)$ instead of the output stream. Empirically, the reduction in filter size through the stream tends to sequentially refine at finer spatial scales. Comparing the enstrophy spectra of the output with and without the output stream at $Re=40$, we see a notable improvement at the largest spatial scale, and a slight improvement at the very small spatial scales. We found that using any smaller filter sizes resulted in an over-smoothing of the reconstructed vorticity fields. This plateau at the high wavenumbers is increasingly absent at higher $Re$, and virtually non-existent at $Re=400$. Instead, the reconstructed fields are overly smoothed at the higher $Re$ values of $200$ and $400$.
Figure 7: The enstrophy spectrum $\Omega(k)$ of the (blue) reconstructed snapshots $\mathcal{A}(\omega)$ and (black) true snapshots $\omega$ at (top left) $Re=40$, (top right) $Re=100$, (bottom left) $Re=200$ and (bottom right) $Re=400$. The red line indicates the enstrophy spectrum at $Re=40$ of a model trained with a single linear convolutional output layer of filter size $(8,8)$.